OpenStudy (evanhelps):

A couch of mass 210kg is being pushed across the floor with an applied force of 4100N. The coefficient of connect friction between the couch and the floor is 0.38. Find the acceleration of the couch.

OpenStudy (evanhelps):

Since it's flat, normal force (Fn) would be equal to mg, which is 2100N. Since Ff=mu*Fn, Ff=798N. So the net force would be 4100N-798N which equals 3302N. Since f=ma, you can solve for a by f/m so a= 15.7 m/s^2 right?

OpenStudy (evanhelps):

Just making sure because I think the answer key I have is wrong and my answer is right.

OpenStudy (evanhelps):

kinetic* btw

OpenStudy (anonymous):

seems right to me.

OpenStudy (anonymous):

maybe they are using 9.8 for gravitational pull instead of the 10 you are using.

jimthompson5910 (jim_thompson5910):

yeah idk how you got g = 10 when g = 9.81 roughly

OpenStudy (evanhelps):

My teacher has us use 10 for the sake of easy numbers

OpenStudy (evanhelps):

Thanks though

OpenStudy (evanhelps):

I tried using 9.8 and the answer was still way off. Their answer was 9.5

jimthompson5910 (jim_thompson5910):

I think you should do this 4100N - 2100N = 2000 N Divide 2000 N by the mass 210 kg to get 2000/210 = 9.5238

jimthompson5910 (jim_thompson5910):

I'd get a second opinion since I'm not the best at physics

OpenStudy (evanhelps):

Well 2100 is the y component of mg and fn, but the object is stationary in the y direction.

OpenStudy (evanhelps):

The acceleration is in the x direction

jimthompson5910 (jim_thompson5910):

hmm then I'm not sure

OpenStudy (evanhelps):

So it's the force exerted minus the force due to friction. And then since f=ma, you can divide the net force by m and get the acceleration.

OpenStudy (evanhelps):

Maybe that's the mistake that they made though

jimthompson5910 (jim_thompson5910):

that is possible

OpenStudy (evanhelps):

Yeah it's just weird for the answer key to be wrong so I posted to see if I was forgetting something

OpenStudy (anonymous):

One thing I learned over the years is to not always trust the answer in the back of a book. With that said, the way you did the problem was correct and the correct answer is \(a\approx 15.8\,m/s^2\) (using \(g=9.8\,m/s^2\). They just forgot to multiply the normal force by the coefficient of friction when computing the net force acting on the couch.